OpenStudy (sand-lock53):

Will medal & fan!!! Plz Help! Algebra - 7 questions!

OpenStudy (sand-lock53):

1. What is the equation of the line with m = 6 that goes through the point (1, 4)? A. y – 4 = 6(x + 1) B. y – 4 = 6(x – 1) C. y + 4 = 6(x – 1) D. y + 4 = 6(x + 1) 2. What is the equation of a horizontal line that has a y-intercept of 3.5? A. y = 3.5 B. y = −3.5 C. x = 3.5 D. y = x− 3.5 These are the 1st two problems

OpenStudy (sand-lock53):

3. What is the equation of the line with m = 2.5 and b = 7? A. y = −2.5x − 7 B. y= 2.5x − 7 C. y = −2.5x + 7 D. y = 2.5x + 7 4. What is the equation of the horizontal line that has a y-intercept of −4? A. y = 4 B. y = −4 C. y=1/4 D. y=-1/4 5. What is the equation of the line that has an x-intercept of −5? A. x = −5 B. x = 5 C. y = x −5 D. y = −x + 5 6. What is the equation of a line that has a y-intercept of 4.5? A. y = 4.5 B. y = −4.5 C. x= 4.5 D. y= x − 4.5 What is the equation of the line from the table of values? A. y = x + 2 B. x= −y − 2 C. y= x − 2 D. x = y + 2

OpenStudy (victoriasushchik):

1Because we have coordinates of a point on the line and slope, let's use Point-Slope Form of linear equation y - y1 = m(x -x1) y - (- 4) = (1/2)(x - 6) y + 4 = (1/2)x - 3 y - (1/2)x = - 3 - 4 Let's multiply both sides of equation by (- 2) x - 2y = 14 ~~~~~~~~~ 6 - 2*(- 4) = 14 6 + 8 = 14 14 = 14

OpenStudy (sand-lock53):

I'm still confused, tho

OpenStudy (sand-lock53):

Do I have to draw a graph?

OpenStudy (victoriasushchik):

@sand-locks53 yes you do